OpenStudy (anonymous):

(3.06 MC) Rachel moved into a new apartment m months ago and paid a $1415 deposit. She pays a monthly rent of $750. How many months has she lived in her apartment if she has paid $8,165?

OpenStudy (whpalmer4):

amount paid = deposit + monthly rent * number of months 8165 = 1415 + 750*m Solve for m to get number of months

OpenStudy (anonymous):

did you get 5.

OpenStudy (whpalmer4):

nope.

OpenStudy (anonymous):

9

OpenStudy (whpalmer4):

5*750 = 3750, + 1415 doesn't get you to 8165

OpenStudy (whpalmer4):

9 is correct. 8165 = 1415 + 9(750) = 1415 + 6750 = 8165

OpenStudy (anonymous):

you the best.

OpenStudy (whpalmer4):

thanks! you're not so bad yourself ;-)